Dynamic viscosity of oil used for lubrication between a shaft and a sleeve is 0.8Ns/m². The shaft is of d=0.5m and rotates at 190rpm. Calculate the horsepower lost in bearing for a sleeve length of 0.1m. Thickness of oil film is 2mm

1
Expert's answer
2022-02-28T16:14:24-0500

"u=\frac{\pi DN}{60}=3.98~\frac ms,"

"\tau=\mu\frac {du}{dy}=1592~\frac N{m^2},"

"F=\tau \pi Dl=180.05~N,"

"T=F\frac D2=36.01~Nm,"

"P=\frac{2\pi NT}{60}=716.48~W."